# LLM-Coordination: Evaluating and Analyzing Multi-agent Coordination Abilities in Large Language Models

### v2

@@ 4.1.1 Setup
Self-play Baselines: For Overcooked we use Proximal Policy Optimization (Schulman et al., 2017) and Population-Based Training (Jaderberg et al., 2017) as baselines for comparison. These baselines were established by Carroll et al. (2019a). The Hanabi challenge has been extensively studied and solved using MARL methods. We use Bayesian Action Decoder (Bard et al., 2020), Simplified Action Decoder (Hu & Foerster, 2021), and Off-Belief Learning (Hu et al., 2021a) as baselines.
@@
Cross-play Baselines: For Overcooked, we use a Behavior Cloning model trained on human data Carroll et al. (2019a) and a Proximal Policy Optimization (PPO) agent trained with the Human Behavior Cloning agent Carroll et al. (2019a) as baselines for comparison. We also use human proxies based on behavior cloning as unseen partners. For Hanabi, we use the Simplified Action Decoder (SAD) as a baseline. We pair our agents with the Off-Belief Learning (Hu et al., 2021a), which was trained to generate grounded policies and adapt to unseen partner agents.
@@
Metrics: We measure the total score achieved by agents in Overcooked, where each delivery provides 20 points to both agents. In the case of Hanabi, the metric is the total number of cards that have been correctly arranged by the players.
4.1.2 Results and Analysis
Overcooked Layouts | |||||
@@ 4.1.2 Results and Analysis
PPO (Schulman et al., 2017) | |||||
PBT (Jaderberg et al., 2017) | |||||
CACGPT-4-turbo | |||||
CACGPT-3.5-turbo | |||||
CACMixtral8x7B |
Class | Method | Score |
RL | BAD (Foerster et al., 2019) | |
SAD (Hu & Foerster, 2021) | ||
OBL (Hu et al., 2021a) | ||
CAC | GPT-4-turbo | |
GPT-3.5-turbo | ||
Mixtral-8x7b |
Method | Score |
LLM+Self verif.+ToM | |
LLM+Self Verif. | |
LLM |
LLM Agents outperform or match state-of-the-art RL methods in coordination games that depend more on understanding the environment.
We observed that LLM agents (w. GPT-4-turbo) outperform or match the overall performance of RL methods across all layouts of Overcooked-AI. Table 1 presents the numerical scores attained by different agents when paired with a partner agent of the same type. This implies that LLM agents outdo RL agents that have been trained together through Self-play without any game-specific training or fine-tuning. It is, however, important to note that LLM agents are significantly slower and larger than RL models and are not fit for real-time use yet (latency (seconds) of with Chain-of-thought and without for GPT-4-turbo). Furthermore, other models we tested, GPT-3.5-turbo and Mixtral8x7b, are faster but fall short of the RL baselines. We also see positive results on the CollabCapture and CollabEscape games with CAC agents (w. GPT-4), achieving a 100% success rate. However, other LLMs are unable to crack CollabEscape (see Appendix D).
LLM agents struggle at effective planning when advanced Theory of Mind reasoning is required. In Hanabi Challenge, LLM agents seem to struggle compared to RL methods (see Table 2). Among all LLMs, GPT-4-turbo performs reasonably well, while other LLMs can barely complete the Hanabi games. We attribute this failure to two factors. First, there is little room for errors in Hanabi. Any misplay or mis-clue leads to the loss of a life token. Second, Hanabi requires much more complex Theory of Mind Reasoning compared to the Overcoked-AI environment. Each action requires agents to actively consider their partner’s beliefs, intentions, and how they would react to implicit communication.
In contrast, Overcooked is fully observable, and its action space consists of actions like pick up an onion from onion_dispenser_0 and place onion in cooker_0. Under most scenarios and layouts, LLMs only need to consider the next best steps based on the state of the environment. For example, We conduct an ablation study of removing partner inventory and location, which reveals minimum impact on overall performance (1 less delivery in Cramped Room and Asymmetric Advantage layouts each in 100 timesteps), showing that the primary challenge for LLMs in games like Overcooked is the Environment Comprehension ability.
LLM agents benefit from auxiliary reasoning modules in imperfection information games with low room for errors. Without the support of auxiliary modules, LLM agents seem to bomb (lose all three lives) in every game (see Table 3). To rectify this, our CAC framework incorporates auxiliary LLM-powered modules, including an Explicit Theory of Mind Reasoning (ToM) module and an Answer Verification (AV) module. The answer verification module is a game-changer in its ability to stop LLM hallucinations about environmental facts from causing fatal mistakes, thus reducing the chance of mis-plays. The ToM reasoning LLM delegates the responsibility of interpreting partner clues and understanding partner needs to different LLMs, allowing the primary LLM to focus on synthesizing the available information to plan the next action.
Overcooked Layouts | |||||
Method | CR | AA | Ring | FC | CC |
BC (Carroll et al., 2019a) | | | | | | | | | | |
PPOBC (Schulman et al., 2017) | | | | | | | | | | |
CACGPT-4-turbo111For CAC, we run a single trial from either position due to cost constraints. In Table 1 we have observed that the performance of CAC agents does not vary by more than one delivery. | | | | | | | | | | |
@@
SAD (Hu & Foerster, 2021) | |||
CACGPT-4-turbo |
LLM Agents are robust to unseen partners. We use Overcooked-AI and the Hanabi challenge as testbeds to evaluate the performance of LLM agents when paired with unseen agents. This task is popularly known as Zero Shot Coordination. For experiments in Overcooked-AI, we pair our LLM agents as well as baselines with proxy-human agents. These proxy human agents are behavior cloning agents trained using human data by Carroll et al. (2019b). As shown in Table 4, we discover that LLM agents outperform both Behavior Cloning as well as PPO agents trained with human data.
For experiments in Hanabi, we pair our agents with Off-Belief Learning (OBL) agents (Hu et al., 2021a). OBL is a MARL strategy that generates grounded clues and actions and is the state-of-the-art method for both self-play and cross-play in Hanabi. OBL agents provide observation-grounded clues and collaborate well with humans. Therefore, we use them as unseen partners in our experiments. Table 5 shows that CAC agents score an average of 15.00 points with the OBL-1 agent compared to their self-play scores of 13.66. This indicates no degradation in coordination abilities with a new partner. The baseline RL method, Simplified Action Decoder (SAD) Hu & Foerster (2021), fails critically when paired with unseen OBL agents, even though it excels at self-play (22.00 points) due to self-play training.
MARL agents trained with self-play struggle when paired with unseen partners in common payoff tasks, because they converge to arbitrary policies that only the two partners involved in the self-play training understand (Carroll et al., 2019a; Bard et al., 2020). Since LLM agents haven’t been explicitly trained to play these games, they base their outputs on the provided textual observation and commonsense knowledge learned from pre-training, and thus are much more robust to unseen partners.
4.2 Coordination QA
Setup.
We assess the performance of 6 Families of Large Language Models (LLMs) Jiang et al. (2023; 2024); Touvron et al. (2023); Chiang et al. (2023); OpenAI (2023) across three dimensions: Environment Comprehension (EC), Theory of Mind Reasoning (ToM), and Joint Planning (JP). For each category, LLMs respond to multiple-choice questions (MCQs), with their responses evaluated against ground-truth answers through fuzzy string matching. To account for the variability in LLM responses, we conduct three trials per model. We also report a Random baseline.
@@ 4.2 Coordination QA
Comparative Results of LLMs in Environment Comprehension, ToM Reasoning, vs. Joint Planning. In Figure 3, we see that most LLMs achieve their best results on the Environment Comprehension question. The best performing LLM GPT-4-turbo gets more than 80% Environment Comprehension Questions correct. The overall performance across LLMs drops on the more challenging Theory of Mind reasoning questions, but GPT-4-turbo is still competent, reaching a 54% accuracy. The overall accuracy of LLMs on Joint Planning questions is still significantly weak, with even the best LLM scoring less than 40%, indicating a large room for improvement in LLMs’ ability to perform coordination reasoning. Another cause for concern is that open-source LLMs perform abysmally at Joint Planning, with some models performing worse than a random baseline.
Variables | ||
ToM | 0.813 | 0.389 |
EC | 0.895 | 0.506 |
Impact of Environment Comprehension and ToM Reasoning abilities on Joint Planning. Having defined Joint Planning as the capacity of an agent to select the appropriate subsequent action based on available information, we argue that proficiency in Environment Comprehension and Theory of Mind Reasoning is crucial for adept Joint Planning, and LLMs that do well at these two will do well at JP. Correlation analysis of ToM and EC with JP across the data from Figure 6 reveals that ToM has a moderate positive correlation to JP, whereas EC shows a strong positive correlation.
5 Related Work
Multi-agent Coordination
In Game Theory, Pure Coordination games are situations where the payoff is commonly shared between all agents. In such situations, cooperating is the best strategy. Various benchmarks and games have been used to evaluate Multi-Agent Coordination abilities over the years including Multiparticle Environment Lowe et al. (2017), Overcooked-AI Carroll et al. (2019a), and the Hanabi Challenge Bard et al. (2020). The foundational work by Carroll et al. (2019a) emphasized the significance of incorporating human data for effective human-ai collaboration. Subsequent research on the Overcooked-AI challenge has pivoted towards enabling self-play-trained agents to coordinate seamlessly with humans within this environment. These studies employ various techniques, including self-play with past agent checkpoints (Strouse et al., 2021), centralized population entropy objectives (Zhao et al., 2023), open-ended objectives using graph theory (Li et al., 2023a), policy ensembles with context-aware mechanisms (Lou et al., 2023), and the incorporation of human biases as linear hidden rewards (Yu et al., 2023), to enhance the training and diversity of AI agents in different scenarios. On the Hanabi Challenge much effort has been made to learn grounded policies Hu et al. (2021b; a) over arbitrary conventions. Embodied environments usually set up in household environments have also been recently used to study multi-agent coordination (Puig et al., 2021; Jain et al., 2020; 2019; Gan et al., 2021). The Overwhelming majority of approaches to coordination problems have focused on utilizing and enhancing Reinforcement Learning methods to solve the problems of multi-agent coordination. In this work, we argue that Large Language Models are an alternative approach to these coordination problems as they show emergent reasoning abilities, demonstrate theory-of-mind-like abilities, and do not converge to policies that cause arbitrary joint interactions.
Planning and Reasoning with Large Language Models
Large Language Models (LLMs) have demonstrated remarkable capabilities of reasoning in natural language (OpenAI, 2023; Ouyang et al., 2022; Chiang et al., 2023), achieving state-of-the-art performance across a spectrum of verbal reasoning tasks. It was then shown that LLMs could be augmented with components like memory, tools, perception, and grounding to create agents that could interact with an external environment (the web, simulators, games, etc.) These LLM agents have shown to be capable of solving long-horizon tasks, playing complex games (Wu et al., 2023; Wang et al., 2023) and interacting with simulated embodied environments (Liang et al., 2022; Song et al., 2022). Zhang et al. (2023b) developed a modular agent framework that was capable of cooperating with partner agents in embodied spatial rearrangement problems, demonstrating increased efficiency through coordination. Zhang et al. (2023a) develop a specialized architecture that enables LLMs to play in the Overcooked-AI environment. Li et al. (2023b) evaluate and show emergent collaborative abilities of LLMs in gamified simulations. In contrast to existing works, our work focuses on evaluating language agents in established pure coordination games where coordination is not an optional efficiency enhancer but rather a necessity.
@@ 6 Conclusion
In this study, we evaluated and analyzed the current large language models in terms of their ability to reason and act in pure coordination games. We introduced the LLM-Coordination benchmark with its two tasks: 1. Agentic Coordination and 2. CoordinationQA. These settings allowed us to conduct holistic comparative studies of LLMs as agents as well as dive deeper into the fine-grained aspects of LLMs as coordination reasoners. We juxtaposed LLM agents with existing Multi-agent Reinforcement Learning agents, discussing the conditions in which LLMs thrive and fail. Finally, we discussed the Theory of Mind Reasoning and Environment Comprehension as prerequisites for coordination and evaluated existing LLMs on these two components.
@@ A.2 State Description
The state description begins by detailing the inventories and of Alice and Bob, respectively. Each inventory (where ) can contain one of the following items: {”onion”, ”plate”, ”cooked soup”}. This inventory information is translated into natural language and incorporated into in the format: “I am holding . Bob is holding .” Such information is vital for inferring the likely subsequent actions of the partner agent.
@@
The state description also incorporates information about the cooker, which is central to the gameplay strategy. Specifically, for each cooker , includes the number of onions currently in the pot. Additionally, provides the operational state of the cooker, denoted as , which can be either ”Off” or ”On”. Lastly, the current condition of the soup in the cooker is represented by , which can take one of the following values: ”Cooking”, ”Cooked”, or ”Not Started”. Thus, the information for cooker is formatted as: “ has onions. is . Soup in is .”
@@ B.1 Game Description
We structure the game description of Hanabi into the overall objective, the rules of the game, and list of conventions based on the H-group conventions han (2024). We do not use advanced conventions like Chop Cards but stick to basic conventions about the card layout, play clues, and save clues.
@@ B.2 State Description
Appendix C Examples of prompts of LLMs used in CAC framework
C.1 Action Generator
The action generator generates a brief explanation and selected the next action to be played.
C.2 Theory of Mind Reasoning LLM
The Theory of Mind Reasoning LLM interprets partner actions and clarifies partner requirements.
C.3 Self Verification LLM
The prompt for the Self Verification LLM is provided as a system prompt. We observed that LLMs were more likely to act as strict verifiers when the prompt is provided as a system prompt.
Appendix D Results of Different LLMs on CollabCapture and CollabEscape
Table 7 summarizes the performance of the three LLMs GPT-4-turbo, GPT-3.5-turbo and Mixtral-8x7b on CollabGames - Collab Capture and Collab Escape. GPT-4-turbo achieves a 100% success rate in both games. In general, CollabEscape was a more difficult game to solve than CollabGames since agents needed to perform sacrificial moves and lure the killer away, requiring more advanced Theory of Mind inference.
Collab Capture | Collab Escape | |||
CAC model | Capture Rate | Turns to Capture | Escape Rate | Turns to Escape |
GPT-4-turbo | 1.00 | 3.99 | 1.00 | 24.67 |
GPT-3.5-turbo | 0.50 | 8.49 | 0.00 | N.A. |
Mixtral-8x7b | 0.75 | 3.88 | 0.00 | N.A. |
Greedy Baseline | 0.50 | 6.00 | 0.00 | N.A. |
Appendix E Generating Questions for CoordinationQA
E.1 Environment Comprehension Questions
The Environment Comprehension (EC) questions are indirect formulations regarding spatial aspects of the layout. In order for an agent to correctly answer an EC question, they must have an understanding of the dynamic details of the current state, the rules of the game, and exhibit spatial awareness. As such, when creating the EC questions, we carefully comb through a given scenario in search of salient points to probe an agent’s understanding of the given environment. Some examples include:
E.2 Theory of Mind Reasoning Questions

### v3

@@ 4.1.2 MARL Agents
Self-play MARL Baselines: For Overcooked we use Proximal Policy Optimization (Schulman et al., 2017) and Population-Based Training (Jaderberg et al., 2017) as baselines for comparison. These baselines were established by Carroll et al. (2019a).
For the Hanabi challenge, we use Bayesian Action Decoder (BAD) (Bard et al., 2020), Simplified Action Decoder (SAD) (Hu and Foerster, 2021), and Off-Belief Learning (Hu et al., 2021a) as MARL baselines for Hanabi. All three baselines achieve near-perfect performance in Self-play.
@@
Cross-play MARL Baselines: For Overcooked, we use a Behavior Cloning model trained on human data Carroll et al. (2019a) and a Proximal Policy Optimization (PPO) agent trained with the Human Behavior Cloning agent Carroll et al. (2019a) as baselines for comparison. We also report Hidden-Utility Self-play (HSP) Yu et al. (2023) as a baseline. We use human proxies based on behavior cloning as unseen partners.
For Hanabi, we use the Simplified Action Decoder (SAD), which is trained through self-play as a baseline. We pair our agents with Off-Belief Learning (Hu et al., 2021a), which was trained to generate grounded policies and adapt to unseen partner agents.
Metrics.
@@
We measure the total score achieved by agents in Overcooked, where each delivery provides 20 points to both agents. In the case of Hanabi, the metric is the total number of cards that have been correctly arranged by the players. For CollabEscape and CollabCapture, we report the success rate of escape or capture across multiple trials and the average turns to capture or escape.
4.2 CoordinationQA
We assess the performance of 5 Families of Large Language Models (LLMs) Jiang et al. (2023, 2024); Touvron et al. (2023); Chiang et al. (2023); OpenAI (2023) across three dimensions: Environment Comprehension (EC), Theory of Mind Reasoning (ToM), and Joint Planning (JP). For each category, LLMs respond to multiple-choice questions (MCQs), with their responses evaluated against ground-truth answers through fuzzy string matching. To account for the variability in LLM responses, we conduct three trials per model. All models being tested are shown the same prompts. We also report a Random baseline.
5 Discussion
Zero-shot LLM Agents match or surpass trained RL methods in Environment-focused Coordination Problems.
We observed that LLM agents (w. GPT-4-turbo) outperform or match the overall performance of RL methods across all layouts of Overcooked-AI. Table 1 presents the numerical scores attained by different agents when paired with a partner agent of the same type. This implies that LLM agents match RL agents that have been explicitly trained through Self-play without any game-specific training or fine-tuning. However, it is important to note that LLM agents are significantly slower and larger than RL models, making them unsuitable for real-time use at present. We also see positive results on the CollabCapture and CollabEscape games, with most LLMs being able to complete both challenges (see Table 2).
| Agent | Score |
| Bayesian Action Decoder | |
| Simplified Action Decoder | |
| Off-Belief Learning | |
| GPT-4-turbo | |
| - (w.o ToM Reasoning) | |
| - (w.o ToM Reasoning & Verif.) | |
| GPT-4o | |
| GPT-3.5-turbo | |
| Mixtral-8x7b |
LLM agents struggle at effective planning when advanced Theory of Mind reasoning is required.
In Hanabi Challenge, LLM agents seem to struggle compared to RL methods (see Table 3). GPT-4-turbo performs reasonably well, while other LLMs can barely complete the games. We attribute this failure to two factors. First, there is little room for errors in Hanabi. Any misplay leads to the loss of a life token. Second, Hanabi requires more complex Theory of Mind Reasoning compared to the Overcooked-AI environment. Each action requires agents to actively consider their partner’s beliefs, intentions, and how they would react to implicit communication. In contrast, Overcooked is fully observable, and its action space consists of actions like pick up an onion from onion_dispenser_0 and place onion in cooker_0. Under most scenarios and layouts, LLMs only need to consider the next best steps based on the state of the environment.
We use correlation study to provide additional validation to these findings. We calculate the Pearson Correlation Coefficient () of the performance of the four LLMs (GPT-4-turbo, GPT-4o, GPT-3.5-turbo, and Mixtral 8x7b) on Agentic Coordination setup (Average score per game) vs. the score on CoordinationQA task. Figure 2 shows a high correlation between Environment Comprehension capabilities and Success at Overcooked, but a more moderate correlation between Theory of Mind Reasoning capabilities and Success at Overcooked. Conversely, in Hanabi, high success is strongly correlated with both Environment Comprehension and Theory of Mind Reasoning Abilities. In CollabEscape we see a higher correlation with ToM reasoning abilities compared to CollabCapture. This correlation study also connects and establishes a positive alignment between LLM-agent performance on the multi-turn agentic task and the single-turn CoordinationQA.
@@ 5 Discussion
| BC | | | | | | | | | | |
| PPOBC | | | | | | | | | | |
| HSP | - | | | | | - | | |
| GPT-4-turbo111For GPT-4-turbo, we run a single trial from either position due to cost and time constraints. | | | | | | | | | | |
@@
| SAD | |||
| GPT-4-turbo |
Auxiliary reasoning strategies like Verification and ToM reasoning help LLMs reason for coordination.
Adding an Answer Verification step significantly reduces fatal mistakes (wrong card plays) caused by LLM hallucinations. Without the support of the Verification step, LLM agents bomb (lose all three lives) before the end of the game in every trial. The ToM reasoning step separates the tasks of interpreting partner clues and generating actions, allowing the LLM to better synthesize available information for action planning. Table 3 shows the impact of ablating the verification and ToM reasoning steps from the LLM Agent. The ToM reasoning step is also useful in the CollabEscape (see table 2) game, as players need to actively consider what their partner needs and act sacrificially if needed. In CollabCapture, it shows a relatively low benefit since agents can observe the positions of all agents as well as doors on the map and infer the correct action based on this environmental context.
Comparative Results of LLMs in Environment Comprehension, ToM Reasoning, and Joint Planning.
@@
In Figure 3, we see that most LLMs achieve their best results on the Environment Comprehension question. The best performing LLM GPT-4-turbo gets more than 80% Environment Comprehension Questions correct. The overall performance across LLMs drops on the more challenging Theory of Mind reasoning questions. Both GPT-4-turbo and GPT-4o do well on the Theory of Mind reasoning questions. The overall accuracy of LLMs on Joint Planning questions is still significantly weak, with even the best LLM scoring less than 40%, indicating a large room for improvement in LLMs’ ability to perform coordination reasoning. Another cause for concern is that open-source LLMs perform abysmally at Joint Planning, with some models performing worse than random.
LLM Agents are robust to unseen partners.
We use Overcooked-AI and the Hanabi challenge as testbeds to evaluate the performance of LLM agents when paired with unseen agents. This task is popularly known as Zero Shot Coordination. In Overcooked-AI, we pair our LLM agents as well as baselines with proxy-human agents. These proxy human agents are behavior cloning agents trained using human data by Carroll et al. (2019b). As shown in Table 4, we discover that LLM agents outperform both Behavior Cloning as well as PPO agents trained with human data. Apart from the Asymmetric Advantages layout, they also match or outperform the Hidden Utility Self-play (HSP) baseline, which is designed to excel at ZSC.
In Hanabi, we pair our agents with Off-Belief Learning (OBL) agents (Hu et al., 2021a). OBL is a MARL strategy that generates grounded clues and actions and is the state-of-the-art method for cross-play in Hanabi. OBL agents provide observation-grounded clues and collaborate well with humans. Therefore, we use them as unseen partners in our experiments. Table 5 shows that the GPT-4-turbo agent scores an average of 15 points with the OBL-1 agent compared to their self-play scores of 13.66, indicating no degradation in performance with a new partner. The baseline RL method, Simplified Action Decoder (SAD) Hu and Foerster (2021), fails critically when paired with unseen OBL agents, even though it excels at self-play (22.00 points) due to self-play training.
@@ 6 Conclusion
In this study, we evaluated and analyzed the current large language models in the context of pure coordination games. We introduced the LLM-Coordination benchmark with its two tasks: 1. Agentic Coordination and 2. CoordinationQA. These settings allowed us to conduct holistic comparative studies of LLMs as agents and dive deeper into the fine-grained aspects of LLMs as coordination reasoners. We juxtaposed LLM agents with existing Multi-agent Reinforcement Learning agents, discussing the conditions in which LLMs thrive and fail. Finally, we discussed the Theory of Mind Reasoning, Environment Comprehension, and Joint Planning as prerequisites for coordination and evaluated existing LLMs on these components.
7 Acknowledgements
This research project has benefitted from the Microsoft Accelerate Foundation Models Research (AFMR) grant program.
8 Limitations
Latency and Compute Requirements: As highlighted in Section 5, effective reasoning for coordination is achievable primarily with larger LLMs like GPT-4-turbo. However, these models are associated with significant latency and require substantial computational resources, making them less suitable for real-time applications where rapid decision-making is crucial.
Initial Prompt Configuration: Achieving optimal reasoning performance necessitates careful manual configuration of the initial prompts that describe the game (Procedural Memory). While this prompt could be extracted from game manuals or existing resources, it still needs to be formatted and designed with the LLM agent in mind. Furthermore, the results for individual games could be improved by letting the LLM generate more text and engineering the prompt. However, we leave these optimizations to future works focused on performance improvement rather than benchmarking.
Manual Curation of Edge Cases: The CoordinationQA suite involves manually curating unambiguous edge cases in coordination games to construct the dataset. This can hinder the ability to scale the benchmark to accommodate new scenarios. Yet, it is important to curate these examples for more reliable studies.
@@ Appendix A Overcooked Implementation Details
In the game Overcooked-AI, two chefs must coordinate their actions to cook and deliver onion soups. Each soup requires three onions as ingredients, which can be found in onion dispensers. The players must add three onions to a cooker to start cooking. Once the three onions are added, the soup starts cooking automatically and requires 20 time steps to complete. Once the soup is cooked, a player needs to pick up a plate, load the soup, and deliver it to the required delivery area. Players must coordinate their actions to effectively deliver soup. Depending on the layouts (see figure 4), players need to adapt to cramped spaces, understand environment layouts for role assignments, pass objects to each other, make way for one another, and find the most effective paths.
@@ A.2 State Description
The state description begins by detailing the inventories and of Alice and Bob, respectively. Each inventory (where ) can contain one of the following items: {"onion", "plate", "cooked soup"}. This inventory information is translated into natural language and incorporated into in the format: “I am holding . Bob is holding .” Such information is vital for inferring the likely subsequent actions of the partner agent.
@@
The state description also incorporates information about the cooker, which is central to the gameplay strategy. Specifically, for each cooker , includes the number of onions currently in the pot. Additionally, provides the operational state of the cooker, denoted as , which can be either "Off" or "On". Lastly, the current condition of the soup in the cooker is represented by , which can take one of the following values: "Cooking", "Cooked", or "Not Started". Thus, the information for cooker is formatted as: “ has onions. is . Soup in is .”
@@ Appendix B Hanabi Implementation Details
Hanabi is a card game in which all players are on the same team. The deck is made up of cards numbered 1 through 5, further divided into five different colors. Players are working together to create these numbered sequences (1 through 5) for each of the five colors. Each card played provides 1 point, and the goal is to obtain all 25 points (5 colors with 5 cards each) by completing the sequence for each color. The three actions a player can take include playing a card, discarding a card, and giving a hint (reveal) to a teammate. The challenge is that players cannot see their own cards. They may, however, see the cards that are in the hands of their teammate(s). This is where hints (reveals) come into play. Players are able to figure out which cards should be played through the reveal mechanism. You can hint (reveal) a teammate about a color or number in their hand, and this will point out to them all cards in their hand that have this color or number. The team starts with 8 reveal tokens but can recover used tokens when discarding a card or upon the completion of a stack. Playing a card carries a risk because the team loses completely if three incorrect cards are played. Further, an incorrectly played card gets sent to the discard pile. Each number of a particular color only has so many copies, so if, for example, the last copy of a green 3 is lost, then the team loses out on not only playing the green 3 but also the green 4 and 5. This is the risk of playing or discarding. With these risks, and since reveal tokens are scarce, players ideally try to make assumptions about implicit information that a reveal may offer, and, ideally, the team converges to particular conventions. Implicit communication, collaboration, and memory are key.
@@ B.1 Game Description
We structure the game description of Hanabi into the overall objective, and the rules of the game.
@@ B.2 State Description
Appendix C Examples of prompts of LLMs used in the Agent framework
C.1 Hanabi LLM Prompt
C.2 Prompt for Theory of Mind Reasoning step
C.3 Prompt for Answer Verification Step
C.4 Ablation on Theory of Mind Reasoning and Verification Steps
| Method | Score | Bomb Rate |
| ToM + GPT-4-turbo + V | ||
| GPT-4-turbo + V | ||
| GPT-4-turbo |
Appendix D Generating Questions for CoordinationQA
D.1 Environment Comprehension Questions
The Environment Comprehension (EC) questions are indirect formulations regarding spatial aspects of the layout. In order for an agent to correctly answer an EC question, they must have an understanding of the dynamic details of the current state and the rules of the game and exhibit spatial awareness. As such, when creating the EC questions, we carefully comb through a given scenario in search of salient points to probe an agent’s understanding of the given environment. Some examples include:
D.2 Theory of Mind Reasoning Questions